\begin{thebibliography}{}

\bibitem[\protect\citename{Aitchison \& Lauder, }1985]{doi:10.2307/2347365}
Aitchison, J., \& Lauder, I.~J. 1985.
\newblock Kernel Density Estimation for Compositional Data.
\newblock {\em Journal of the Royal Statistical Society. Series C (Applied
  Statistics)}, {\bf 34}(2), 129--137.
\newblock \href{https://doi.org/10.2307/2347365}{doi:10.2307/2347365}.

\bibitem[\protect\citename{Arenbaev, }1976]{MR0478288}
Arenbaev, N.~K. 1976.
\newblock Asymptotic behavior of the multinomial distribution.
\newblock {\em Teor. Veroyatnost. i Primenen.}, {\bf 21}(4), 826--831.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0478288}{MR0478288}.

\bibitem[\protect\citename{Babu {\em et~al.}, }2002]{MR1910059}
Babu, G.~J., Canty, A.~J., \& Chaubey, Y.~P. 2002.
\newblock Application of {B}ernstein polynomials for smooth estimation of a
  distribution and density function.
\newblock {\em J. Statist. Plann. Inference}, {\bf 105}(2), 377--392.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1910059}{MR1910059}.

\bibitem[\protect\citename{Belalia {\em et~al.}, }2017]{MR3630225}
Belalia, M., Bouezmarni, T., \& Leblanc, A. 2017.
\newblock Smooth conditional distribution estimators using {B}ernstein
  polynomials.
\newblock {\em Comput. Statist. Data Anal.}, {\bf 111}, 166--182.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3630225}{MR3630225}.

\bibitem[\protect\citename{Belalia {\em et~al.}, }2019]{MR3983257}
Belalia, M., Bouezmarni, T., \& Leblanc, A. 2019.
\newblock Bernstein conditional density estimation with application to
  conditional distribution and regression functions.
\newblock {\em J. Korean Statist. Soc.}, {\bf 48}(3), 356--383.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3983257}{MR3983257}.

\bibitem[\protect\citename{Bouezmarni \& Rolin, }2003]{MR1985506}
Bouezmarni, T., \& Rolin, J.-M. 2003.
\newblock Consistency of the beta kernel density function estimator.
\newblock {\em Canad. J. Statist.}, {\bf 31}(1), 89--98.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1985506}{MR1985506}.

\bibitem[\protect\citename{Bouezmarni \& Rolin, }2007]{MR2351744}
Bouezmarni, T., \& Rolin, J.-M. 2007.
\newblock Bernstein estimator for unbounded density function.
\newblock {\em J. Nonparametr. Stat.}, {\bf 19}(3), 145--161.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2351744}{MR2351744}.

\bibitem[\protect\citename{Bouezmarni {\em et~al.}, }2007]{MR2395599}
Bouezmarni, T., M., Mesfioui, \& Rolin, J.~M. 2007.
\newblock $L_1$-rate of convergence of smoothed histogram.
\newblock {\em Statist. Probab. Lett.}, {\bf 77}(14), 1497--1504.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2395599}{MR2395599}.

\bibitem[\protect\citename{Brown \& Chen, }1999]{MR1685301}
Brown, B.~M., \& Chen, S.~X. 1999.
\newblock Beta-{B}ernstein smoothing for regression curves with compact
  support.
\newblock {\em Scand. J. Statist.}, {\bf 26}(1), 47--59.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1685301}{MR1685301}.

\bibitem[\protect\citename{Chen, }1999]{MR1718494}
Chen, S.~X. 1999.
\newblock Beta kernel estimators for density functions.
\newblock {\em Comput. Statist. Data Anal.}, {\bf 31}(2), 131--145.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1718494}{MR1718494}.

\bibitem[\protect\citename{Chen, }2000]{MR1742101}
Chen, S.~X. 2000.
\newblock Beta kernel smoothers for regression curves.
\newblock {\em Statist. Sinica}, {\bf 10}(1), 73--91.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1742101}{MR1742101}.

\bibitem[\protect\citename{Curtis \& Ghosh, }2011]{MR2782409}
Curtis, S.~M., \& Ghosh, S.~K. 2011.
\newblock A variable selection approach to monotonic regression with
  {B}ernstein polynomials.
\newblock {\em J. Appl. Stat.}, {\bf 38}(5), 961--976.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2782409}{MR2782409}.

\bibitem[\protect\citename{Gawronski, }1985]{MR0791719}
Gawronski, W. 1985.
\newblock Strong laws for density estimators of {B}ernstein type.
\newblock {\em Period. Math. Hungar}, {\bf 16}(1), 23--43.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0791719}{MR0791719}.

\bibitem[\protect\citename{Gawronski \& Stadtm\"uller, }1981]{MR0638651}
Gawronski, W., \& Stadtm\"uller, U. 1981.
\newblock Smoothing histograms by means of lattice and continuous
  distributions.
\newblock {\em Metrika}, {\bf 28}(3), 155--164.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0638651}{MR0638651}.

\bibitem[\protect\citename{Ghosal, }2001]{MR1873330}
Ghosal, S. 2001.
\newblock Convergence rates for density estimation with {B}ernstein
  polynomials.
\newblock {\em Ann. Statist.}, {\bf 29}(5), 1264--1280.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1873330}{MR1873330}.

\bibitem[\protect\citename{Graham {\em et~al.}, }1994]{MR1397498}
Graham, R.~L., Knuth, D.~E., \& Patashnik, O. 1994.
\newblock {\em Concrete mathematics}. Second edn.
\newblock Addison-Wesley Publishing Company, Reading, MA.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1397498}{MR1397498}.

\bibitem[\protect\citename{Guan, }2016]{MR3488598}
Guan, Z. 2016.
\newblock Efficient and robust density estimation using {B}ernstein type
  polynomials.
\newblock {\em J. Nonparametr. Stat.}, {\bf 28}(2), 250--271.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3488598}{MR3488598}.

\bibitem[\protect\citename{Guan, }2017]{MR3740722}
Guan, Z. 2017.
\newblock Bernstein polynomial model for grouped continuous data.
\newblock {\em J. Nonparametr. Stat.}, {\bf 29}(4), 831--848.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3740722}{MR3740722}.

\bibitem[\protect\citename{Igarashi \& Kakizawa, }2014]{MR3174309}
Igarashi, G., \& Kakizawa, Y. 2014.
\newblock On improving convergence rate of {B}ernstein polynomial density
  estimator.
\newblock {\em J. Nonparametr. Stat.}, {\bf 26}(1), 61--84.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3174309}{MR3174309}.

\bibitem[\protect\citename{Johnson {\em et~al.}, }1997]{MR1429617}
Johnson, N.~L., Kotz, S., \& Balakrishnan, N. 1997.
\newblock {\em Discrete multivariate distributions}.
\newblock Wiley Series in Probability and Statistics: Applied Probability and
  Statistics.
\newblock John Wiley \& Sons, Inc., New York.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1429617}{MR1429617}.

\bibitem[\protect\citename{Kakizawa, }2004]{MR2068610}
Kakizawa, Y. 2004.
\newblock Bernstein polynomial probability density estimation.
\newblock {\em J. Nonparametr. Stat.}, {\bf 16}(5), 709--729.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2068610}{MR2068610}.

\bibitem[\protect\citename{Leblanc, }2009]{MR2488150}
Leblanc, A. 2009.
\newblock Chung-{S}mirnov property for {B}ernstein estimators of distribution
  functions.
\newblock {\em J. Nonparametr. Stat.}, {\bf 21}(2), 133--142.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2488150}{MR2488150}.

\bibitem[\protect\citename{Leblanc, }2010]{MR2662607}
Leblanc, A. 2010.
\newblock A bias-reduced approach to density estimation using {B}ernstein
  polynomials.
\newblock {\em J. Nonparametr. Stat.}, {\bf 22}(3-4), 459--475.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2662607}{MR2662607}.

\bibitem[\protect\citename{Leblanc, }2012a]{MR2960952}
Leblanc, A. 2012a.
\newblock On estimating distribution functions using {B}ernstein polynomials.
\newblock {\em Ann. Inst. Statist. Math.}, {\bf 64}(5), 919--943.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2960952}{MR2960952}.

\bibitem[\protect\citename{Leblanc, }2012b]{MR2925964}
Leblanc, A. 2012b.
\newblock On the boundary properties of {B}ernstein polynomial estimators of
  density and distribution functions.
\newblock {\em J. Statist. Plann. Inference}, {\bf 142}(10), 2762--2778.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR2925964}{MR2925964}.

\bibitem[\protect\citename{Lu, }2015]{MR3412755}
Lu, L. 2015.
\newblock On the uniform consistency of the {B}ernstein density estimator.
\newblock {\em Statist. Probab. Lett.}, {\bf 107}, 52--61.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3412755}{MR3412755}.

\bibitem[\protect\citename{Mosimann, }1962]{MR143299}
Mosimann, J.~E. 1962.
\newblock On the compound multinomial distribution, the multivariate {$\beta
  $}-distribution, and correlations among proportions.
\newblock {\em Biometrika}, {\bf 49}, 65--82.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR143299}{MR143299}.

\bibitem[\protect\citename{Newcomer, }2008]{Newcomer2008phd}
Newcomer, J.~T. 2008.
\newblock {\em Estimation procedures for multinomial models with
  overdispersion}.
\newblock Ph{D} thesis, University of Maryland.

\bibitem[\protect\citename{Newcomer {\em et~al.},
  }2008]{Newcomer_et_al_2008_tech_report}
Newcomer, J.~T., Neerchal, N.~K., \& Morel, J.~G. 2008.
\newblock Computation of higher order moments from two multinomial
  overdispersion likelihood models.
\newblock {\em Preprint},  1--11.
\newblock \newline [URL]
  ~\url{http://www.math.umbc.edu/~kogan/technical_papers/2008/Newcomer_Nagaraj_Morel.pdf}.

\bibitem[\protect\citename{Ouimet, }2018]{MR3825458}
Ouimet, F. 2018.
\newblock Complete monotonicity of multinomial probabilities and its
  application to {B}ernstein estimators on the simplex.
\newblock {\em J. Math. Anal. Appl.}, {\bf 466}(2), 1609--1617.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3825458}{MR3825458}.

\bibitem[\protect\citename{Ouimet, }2019]{Ouimet2019phd}
Ouimet, F. 2019.
\newblock {\em Extremes of log-correlated random fields and the {R}iemann-zeta
  function, and some asymptotic results for various estimators in statistics}.
\newblock Ph{D} thesis, Universit\'e de Montr\'eal.
\newblock \url{http://hdl.handle.net/1866/22667}.

\bibitem[\protect\citename{Ouimet, }2020a]{arXiv:2002.07758}
Ouimet, F. 2020a.
\newblock Asymptotic properties of {B}ernstein estimators on the simplex.
\newblock {\em Preprint},  1--27.
\newblock \href{https://arxiv.org/abs/2002.07758}{arXiv:2002.07758}.

\bibitem[\protect\citename{Ouimet, }2020b]{arXiv:2006.11756}
Ouimet, F. 2020b.
\newblock Asymptotic properties of {B}ernstein estimators on the simplex. Part
  2: the boundary case.
\newblock {\em Preprint},  1--23.
\newblock \href{https://arxiv.org/abs/2006.11756}{arXiv:2006.11756}.

\bibitem[\protect\citename{Ouimet, }2020c]{arXiv:2002.06956}
Ouimet, F. 2020c.
\newblock Density estimation using Dirichlet kernels.
\newblock {\em Preprint},  1--39.
\newblock \href{https://arxiv.org/abs/2002.06956}{arXiv:2002.06956}.

\bibitem[\protect\citename{Ouimet, }2020d]{arXiv:2006.09059}
Ouimet, F. 2020d.
\newblock Explicit formulas for the joint third and fourth central moments of
  the multinomial distribution.
\newblock {\em Preprint},  1--7.
\newblock \href{https://arxiv.org/abs/2006.09059}{arXiv:2006.09059}.

\bibitem[\protect\citename{Ouimet, }2020e]{arXiv:2001.08512}
Ouimet, F. 2020e.
\newblock A precise local limit theorem for the multinomial distribution and
  some applications.
\newblock {\em Preprint},  1--20.
\newblock \href{https://arxiv.org/abs/2001.08512}{arXiv:2001.08512}.

\bibitem[\protect\citename{Petrone, }1999a]{MR1703623}
Petrone, S. 1999a.
\newblock Bayesian density estimation using {B}ernstein polynomials.
\newblock {\em Canad. J. Statist.}, {\bf 27}(1), 105--126.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1703623}{MR1703623}.

\bibitem[\protect\citename{Petrone, }1999b]{MR1712051}
Petrone, S. 1999b.
\newblock Random {B}ernstein polynomials.
\newblock {\em Scand. J. Statist.}, {\bf 26}(3), 373--393.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1712051}{MR1712051}.

\bibitem[\protect\citename{Petrone \& Wasserman, }2002]{MR1881846}
Petrone, S., \& Wasserman, L. 2002.
\newblock Consistency of {B}ernstein polynomial posteriors.
\newblock {\em J. Roy. Statist. Soc. Ser. B}, {\bf 64}(1), 79--100.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1881846}{MR1881846}.

\bibitem[\protect\citename{Stadtm\"uller, }1983]{MR0726014}
Stadtm\"uller, U. 1983.
\newblock Asymptotic distributions of smoothed histograms.
\newblock {\em Metrika}, {\bf 30}(3), 145--158.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0726014}{MR0726014}.

\bibitem[\protect\citename{Stadtm\"uller, }1986]{MR0858109}
Stadtm\"uller, U. 1986.
\newblock Asymptotic properties of nonparametric curve estimates.
\newblock {\em Period. Math. Hungar.}, {\bf 17}(2), 83--108.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0858109}{MR0858109}.

\bibitem[\protect\citename{Tenbusch, }1994]{MR1293514}
Tenbusch, A. 1994.
\newblock Two-dimensional {B}ernstein polynomial density estimators.
\newblock {\em Metrika}, {\bf 41}(3-4), 233--253.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1293514}{MR1293514}.

\bibitem[\protect\citename{Tenbusch, }1997]{MR1437794}
Tenbusch, A. 1997.
\newblock Nonparametric curve estimation with {B}ernstein estimates.
\newblock {\em Metrika}, {\bf 45}(1), 1--30.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR1437794}{MR1437794}.

\bibitem[\protect\citename{Turnbull \& Ghosh, }2014]{MR3139345}
Turnbull, B.~C., \& Ghosh, S.~K. 2014.
\newblock Unimodal density estimation using {B}ernstein polynomials.
\newblock {\em Comput. Statist. Data Anal.}, {\bf 72}, 13--29.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR3139345}{MR3139345}.

\bibitem[\protect\citename{Vitale, }1975]{MR0397977}
Vitale, R.~A. 1975.
\newblock Bernstein polynomial approach to density function estimation.
\newblock {\em Pages  87--99 of:} {\em Statistical {I}nference and {R}elated
  {T}opics}.
\newblock Academic Press, New York.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR0397977}{MR0397977}.

\bibitem[\protect\citename{Wishart, }1949]{MR33996}
Wishart, J. 1949.
\newblock Cumulants of multivariate multinomial distributions.
\newblock {\em Biometrika}, {\bf 36}, 47--58.
\newblock \href{http://www.ams.org/mathscinet-getitem?mr=MR33996}{MR33996}.

\end{thebibliography}

\documentclass[final,3p,10pt,authoryear]{elsarticle}

\makeatletter
\def\ps@pprintTitle{%
 \let\@oddhead\@empty
 \let\@evenhead\@empty
 \def\@oddfoot{}%
 \let\@evenfoot\@oddfoot}
\makeatother

\usepackage{amssymb,amsthm,mathtools}
\usepackage[colorlinks]{hyperref}
\usepackage{natbib}

\DeclareRobustCommand{\brkbinom}{\genfrac\{\}{0pt}{}}
\usepackage{xcolor,dsfont,geometry}
\newtheorem{Theorem}{Theorem}
\newtheorem{Lemma}{Lemma}

\makeatletter \@addtoreset{equation}{section} \makeatother

\newcommand{\N}{\mathbb{N}}
\newcommand{\R}{\mathbb{R}}

\newcommand{\EE}{\mathbb{E}}

\newcommand{\bb}[1]{\boldsymbol{#1}}

\newcommand{\leqdef}{\vcentcolon=}

\newcommand{\blue}{\color{blue}}
\newcommand{\ind}{\mathds{1}}

\allowdisplaybreaks

\begin{document}

\begin{frontmatter}

    \title{General formulas for the central and non-central moments of the multinomial distribution}%

    \author[a1]{Fr\'ed\'eric Ouimet\texorpdfstring{\corref{cor1}\fnref{fn1}}{)}}%

    \address[a1]{California Institute of Technology, Pasadena, USA.}%

    \cortext[cor1]{Corresponding author}%
    \ead{ouimetfr@caltech.edu}%

    \begin{abstract}
        We present the first general formulas for the central and non-central moments of the multinomial distribution, using a combinatorial argument and the factorial moments previously obtained in \texorpdfstring{\cite{MR143299}}{Mosimann (1962)}. We use the formulas to give explicit expressions for all the non-central moments up to order~8 and all the central moments up to order~4. These results expand significantly on those in \texorpdfstring{\cite{Newcomer2008phd,Newcomer_et_al_2008_tech_report}}{Newcomer (2008) and Newcomer et al.\ (2008)}, where the non-central moments were calculated up to order~4.
    \end{abstract}

    \begin{keyword}
        multinomial distribution \sep higher moments \sep central moments \sep non-central moments
        \MSC[2010]{Primary : 62E15 Secondary : 60E05}
    \end{keyword}

\end{frontmatter}

\section{The multinomial distribution}\label{sec:intro}

    For any $d\in \N$, the $d$-dimensional (unit) simplex is defined by $\mathcal{S} \leqdef \big\{\bb{x}\in [0,1]^d : \sum_{i=1}^d x_i \leq 1\big\}$, and the probability mass function $\bb{k}\mapsto P_{\bb{k},m}(\bb{x})$ for $\bb{\xi} \leqdef (\xi_1,\xi_2,\dots,\xi_d) \sim \mathrm{Multinomial}\hspace{0.2mm}(m,\bb{x})$ is defined by
    \begin{equation}\label{eq:multinomial.probability}
        P_{\bb{k},m}(\bb{x}) \leqdef \frac{m!}{(m - \sum_{i=1}^d k_i)! \prod_{i=1}^d k_i!} \cdot (1 - \sum_{i=1}^d x_i)^{m - \sum_{i=1}^d k_i} \prod_{i=1}^d x_i^{k_i}, \quad \bb{k}\in \N_0^d \cap m\mathcal{S},
    \end{equation}
    where $m\in \N$ and $\bb{x}\in \mathcal{S}$.
    If $x_{d+1}\leqdef 1-\sum_{i=1}^d x_i$, then \eqref{eq:multinomial.probability} is just a reparametrization of $(\bb{\xi},1-\sum_{i=1}^d \xi_i)\sim \mathrm{Multinomial}\hspace{0.2mm}(m,(\bb{x},x_{d+1}))$ where $\sum_{i=1}^{d+1} x_i = 1$.
    In this paper, our main goal is to give general formulas for the non-central and central moments of \eqref{eq:multinomial.probability}, namely
    \begin{equation}\label{eq:main.goal}
        \EE\Big[\prod_{i=1}^d \xi_i^{p_i}\Big] \quad \text{and} \quad \EE\Big[\prod_{i=1}^d (\xi_i - \EE[\xi_i])^{p_i}\Big], \qquad p_1,p_2,\dots,p_d\in \N_0.
    \end{equation}
    We obtain the formulas using a combinatorial argument and the general expression for the factorial moments found in \cite{MR143299}, which we register in the lemma below.
    \begin{Lemma}[{\blue Factorial moments}]\label{lem:factorial.moments}
        Let $\bb{\xi}\sim \mathrm{Multinomial}\hspace{0.3mm}(m,\bb{x})$.
        Then, for all $k_1, k_2, \dots, k_d\in \N_0$,
        \begin{equation}\label{eq:factorial.moments}
            \EE\Big[\prod_{i=1}^d \xi_i^{(k_i)}\Big] = m^{(\sum_{i=1}^d k_i)} \prod_{i=1}^d x^{k_i},
        \end{equation}
        where $m^{(k)} \leqdef m (m-1) \dots (m-k+1)$ denotes the {\it $k$-th order falling factorial of $m$}.
    \end{Lemma}
    The formulas that we develop for the expectations in \eqref{eq:main.goal} will be used to compute explicitly all the non-central moments up to order~8 and all the central moments up to order~4, which expands on the third and fourth order non-central moments that were previously calculated in \cite[Appendix A.1]{Newcomer2008phd}.
    We should also mention that explicit formulas for several lower-order (mixed) cumulants were presented in \cite{MR33996} (see also \cite[p.37]{MR1429617}), but not for the moments.

\section{Motivation}\label{sec:motivation}

    To the best of our knowledge, general formulas for the central and non-central moments of the multinomial distribution have never been derived in the literature.
    The central moments can arise naturally, for example, when studying asymptotic properties, via Taylor expansions, of statistical estimators involving the multinomial distribution.
    For a given sequence of i.i.d.\ observations $\bb{X}_1,\bb{X}_2,\dots,\bb{X}_n$, two examples of such estimators are the Bernstein estimator for the cumulative distribution function
    \begin{equation}\label{eq:cdf.Bernstein.estimator}
        F_{n,m}^{\star}(\bb{x}) \leqdef \sum_{\bb{k}\in \N_0^d \cap m \mathcal{S}} \frac{1}{n} \sum_{i=1}^n \ind_{(-\bb{\infty},\frac{\bb{k}}{m}]}(\bb{X}_i) P_{\bb{k},m}(\bb{x}), \quad \bb{x}\in \mathcal{S}, ~~m,n \in \N,
    \end{equation}
    and the Bernstein estimator for the density function (also called smoothed histogram)
    \begin{equation}\label{eq:histogram.estimator}
        \hat{f}_{n,m}(\bb{x}) \leqdef \sum_{\bb{k}\in \N_0^d \cap (m-1) \mathcal{S}} \frac{m^d}{n} \sum_{i=1}^n \ind_{(\frac{\bb{k}}{m}, \frac{\bb{k} + 1}{m}]}(\bb{X}_i) P_{\bb{k},m-1}(\bb{x}), \quad \bb{x}\in \mathcal{S}, ~~m,n \in \N,
    \end{equation}
    over the $d$-dimensional simplex.
    Some of their asymptotic properties were investigated by \cite{MR0397977,MR0638651,MR0726014,MR0791719,MR0858109,MR1437794,MR1703623,MR1712051,MR1873330,MR1881846,MR1910059,MR2068610,MR2351744,MR2395599,MR2488150,MR2662607,MR2782409,MR2960952,MR2925964,MR3174309,MR3139345,MR3412755,MR3488598,MR3740722,MR3630225,MR3983257} when $d = 1$, by \cite{MR1293514} when $d = 2$, and by \cite{arXiv:2002.07758,arXiv:2006.11756} for all $d\geq 1$, using a local limit theorem from \cite{arXiv:2001.08512} for the multinomial distribution (see also \cite{MR0478288}).
    The estimator \eqref{eq:histogram.estimator} is a discrete analogue of the Dirichlet kernel estimator introduced by \cite{doi:10.2307/2347365} and studied theoretically in \cite{MR1685301,MR1718494,MR1742101,MR1985506} when $d = 1$ (among others), and in \cite{arXiv:2002.06956} for all $d\geq 1$.

\section{Results}\label{sec:main.results}

    First, we give a general formula of the non-central moments of the multinomial distribution in \eqref{eq:multinomial.probability}.

    \begin{Theorem}[{\blue Non-central moments}]\label{thm:non.central.moments}
        Let $\bb{\xi}\sim \mathrm{Multinomial}\hspace{0.3mm}(m,\bb{x})$.
        Then, for all $p_1, p_2, \dots, p_d\in \N_0$,
        \begin{equation}
            \EE\Big[\prod_{i=1}^d \xi_i^{p_i}\Big] = \sum_{k_1=0}^{p_1} \dots \sum_{k_d=0}^{p_d} m^{(\sum_{i=1}^d k_i)} \prod_{i=1}^d \brkbinom{p_i}{k_i} x_i^{k_i},
        \end{equation}
        where $\brkbinom{p}{k}$ denotes a Stirling number of the second kind (i.e., the number of ways to partition a set of $p$ objects into $k$ non-empty subsets).
    \end{Theorem}

    \begin{proof}
        We have the following well-known relation between the power $p\in \N_0$ of a number $x\in \R$ and the falling factorials of $x$:
        \begin{equation}
            x^p = \sum_{k=0}^p \brkbinom{p}{k} \, x^{(k)}.
        \end{equation}
        See, e.g., \cite[p.262]{MR1397498}.
        Apply this relation to every $\xi_i^{p_i}$ and use the linearity of the expectation to get
        \begin{equation}
            \EE\Big[\prod_{i=1}^d \xi_i^{p_i}\Big] = \sum_{k_1=0}^{p_1} \dots \sum_{k_d=0}^{p_d} \brkbinom{p_1}{k_1} \dots \brkbinom{p_d}{k_d} \, \EE\Big[\prod_{i=1}^d \xi_i^{(k_i)}\Big],
        \end{equation}
        The conclusion follows from Lemma~\ref{lem:factorial.moments}.
    \end{proof}

    We can now deduce a general formula for the central moments of the multinomial distribution.

    \begin{Theorem}[{\blue Central moments}]\label{thm:central.moments}
        Let $\bb{\xi}\sim \mathrm{Multinomial}\hspace{0.3mm}(m,\bb{x})$.
        Then, for all $p_1, p_2, \dots, p_d\in \N_0$,
        \begin{equation}
            \EE\Big[\prod_{i=1}^d (\xi_i - \EE[\xi_i])^{p_i}\Big] = \sum_{\ell_1=0}^{p_1} \dots \sum_{\ell_d=0}^{p_d} \sum_{k_1=0}^{\ell_1} \dots \sum_{k_d=0}^{\ell_d} m^{(\sum_{i=1}^d k_i)} (-m)^{\sum_{i=1}^d (p_i - \ell_i)} \prod_{i=1}^d \binom{p_i}{\ell_i} \brkbinom{\ell_i}{k_i} x_i^{p_i - \ell_i + k_i},
        \end{equation}
        where $\binom{p}{\ell}$ denotes the binomial coefficient $\frac{p!}{\ell! (p - \ell)!}$.
    \end{Theorem}

    \begin{proof}
        By applying the binomial formula to each factor $(\xi_i - \EE[\xi_i])^{p_i}$ and using the fact that $\EE[\xi_i] = m x_i$ for all $i\in \{1,2,\dots,d\}$, note that
        \begin{equation}
            \EE\Big[\prod_{i=1}^d (\xi_i - \EE[\xi_i])^{p_i}\Big] = \sum_{\ell_1=0}^{p_1} \dots \sum_{\ell_d=0}^{p_d} \EE\Big[\prod_{i=1}^d \xi_i^{\ell_i}\Big] \cdot \prod_{i=1}^d \binom{p_i}{\ell_i} (- m x_i)^{p_i - \ell_i}.
        \end{equation}
        The conclusion follows from Theorem~\ref{thm:non.central.moments}.
    \end{proof}

\section{Numerical implementation}

    The formulas in Theorem~\ref{thm:non.central.moments} and Theorem~\ref{thm:central.moments} can be implemented in \texttt{Mathematica} as follows:
    \begin{verbatim}
    NonCentral[m_, x_, p_, d_] :=
        Sum[FactorialPower[m, Sum[k[i], {i, 1, d}]] *
            Product[StirlingS2[p[[i]], k[i]] * x[[i]] ^ k[i], {i, 1, d}], ##] & @@
                ({k[#], 0, p[[#]]} & /@ Range[d]);
    Central[m_, x_, p_, d_] :=
        Sum[Sum[FactorialPower[m, Sum[k[i], {i, 1, d}]] *
            (-m) ^ Sum[p[[i]] - ell[i], {i, 1, d}] *
                Product[Binomial[p[[i]], ell[i]] * StirlingS2[ell[i], k[i]] *
                    x[[i]] ^ (p[[i]] - ell[i] + k[i]), {i, 1, d}], ##] & @@
                        ({k[#], 0, ell[#]} & /@ Range[d]), ##] & @@
                            ({ell[#], 0, p[[#]]} & /@ Range[d]);
    \end{verbatim}

\section{Explicit formulas}

    In \cite{Newcomer2008phd}, explicit expressions for the non-central moments of order~3 and 4 where obtained for the multinomial distribution, see also \cite{Newcomer_et_al_2008_tech_report,arXiv:2006.09059}.
    To expand on those results, we use the formula from Theorem~\ref{thm:non.central.moments} in the two subsections below to calculate (explicitly) all the non-central moments up to order~8 and all the central moments up to order~4.

    \vspace{3mm}
    Here is a table of the Stirling numbers of the second kind that we will use in our calculations:
    \small
    \begin{equation*}
        \begin{aligned}
            &\brkbinom{0}{0} = 1, \\[0.5mm]
            &\brkbinom{1}{0} = 0, ~ \brkbinom{1}{1} = 1, \\[0.5mm]
            &\brkbinom{2}{0} = 0, ~ \brkbinom{2}{1} = 1, \brkbinom{2}{2} = 1, \\[0.5mm]
            &\brkbinom{3}{0} = 0, ~ \brkbinom{3}{1} = 1, \brkbinom{3}{2} = 3, ~ \brkbinom{3}{3} = 1, \\[0.5mm]
            &\brkbinom{4}{0} = 0, ~ \brkbinom{4}{1} = 1, \brkbinom{4}{2} = 7, ~ \brkbinom{4}{3} = 6, \brkbinom{4}{4} = 1, \\[0.5mm]
            &\brkbinom{5}{0} = 0, ~ \brkbinom{5}{1} = 1, \brkbinom{5}{2} = 15, ~ \brkbinom{5}{3} = 25, \brkbinom{5}{4} = 10, ~ \brkbinom{5}{5} = 1, \\[0.5mm]
            &\brkbinom{6}{0} = 0, ~ \brkbinom{6}{1} = 1, \brkbinom{6}{2} = 31, ~ \brkbinom{6}{3} = 90, \brkbinom{6}{4} = 65, ~ \brkbinom{6}{5} = 15, \brkbinom{6}{6} = 1, \\[0.5mm]
            &\brkbinom{7}{0} = 0, ~ \brkbinom{7}{1} = 1, \brkbinom{7}{2} = 63, ~ \brkbinom{7}{3} = 301, \brkbinom{7}{4} = 350, ~ \brkbinom{7}{5} = 140, \brkbinom{7}{6} = 21, ~ \brkbinom{7}{7} = 1, \\[0.5mm]
            &\brkbinom{8}{0} = 0, ~ \brkbinom{8}{1} = 1, \brkbinom{8}{2} = 127, ~ \brkbinom{8}{3} = 966, \brkbinom{8}{4} = 1701, ~ \brkbinom{8}{5} = 1050, \brkbinom{8}{6} = 266, ~ \brkbinom{8}{7} = 28, ~ \brkbinom{8}{8} = 1.
        \end{aligned}
    \end{equation*}

    \subsection{Computation of the non-central moments up to order~8}

    By applying the general expression in Theorem~\ref{thm:non.central.moments} and by removing the Stirling numbers $\brkbinom{p_i}{k_i}$ that are equal to $0$, we get the following results directly.

    \vspace{3mm}
    \noindent
    \underline{Order $1$:} For any $j_1\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}]
        &= x_{j_1} m.
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $2$:} For any distinct $j_1, j_2\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}^2]
        &= x_{j_1} \big[m + m^{(2)} x_{j_1}\big], \\[1.1mm]
        \EE[\xi_{j_1} \xi_{j_2}]
        &= x_{j_1} x_{j_2} m^{(2)}.
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $3$:} For any distinct $j_1, j_2, j_3\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}^3]
        &= x_{j_1} \big[m + 3 m^{(2)} x_{j_1} + m^{(3)} x_{j_1}^2\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}]
        &= x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} x_{j_1}\big], \\[1.1mm]
        \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} m^{(3)}.
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $4$:} For any distinct $j_1, j_2, j_3, j_4\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}^4]
        &= x_{j_1} \big[m + 7 m^{(2)} x_{j_1} + 6 m^{(3)} x_{j_1}^2 + m^{(4)} x_{j_1}^3\big], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}]
        &= x_{j_1} x_{j_2} \big[m^{(2)} + 3 m^{(3)} x_{j_1} + m^{(4)} x_{j_1}^2\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2]
        &= x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} (x_{j_1} + x_{j_2}) + m^{(4)} x_{j_1} x_{j_2}\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2} \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \big[m^{(3)} + m^{(4)} x_{j_1}\big], \\[1.1mm]
        \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} m^{(4)}.
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $5$:} For any distinct $j_1, j_2, j_3, j_4, j_5\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}^5]
        &= x_{j_1} \big[m + 15 m^{(2)} x_{j_1} + 25 m^{(3)} x_{j_1}^2 + 10 m^{(4)} x_{j_1}^3 + m^{(5)} x_{j_1}^4\big], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}]
        &= x_{j_1} x_{j_2} \big[m^{(2)} + 7 m^{(3)} x_{j_1} + 6 m^{(4)} x_{j_1}^2 + m^{(5)} x_{j_1}^3\big], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^2]
        &= x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} (3 x_{j_1} + x_{j_2}) + m^{(4)} (x_{j_1}^2 + 3 x_{j_1} x_{j_2}) + m^{(5)} x_{j_1}^2 x_{j_2}\big], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2} \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \big[m^{(3)} + 3 m^{(4)} x_{j_1} + m^{(5)} x_{j_1}^2\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \big[m^{(3)} + m^{(4)} (x_{j_1} + x_{j_2}) + m^{(5)} x_{j_1} x_{j_2}\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2} \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \big[m^{(4)} + m^{(5)} x_{j_1}\big], \\[1.1mm]
        \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} m^{(5)}.
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $6$:} For any distinct $j_1, j_2, j_3, j_4, j_5, j_6\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}^6]
        &= x_{j_1} \big[m + 31 m^{(2)} x_{j_1} + 90 m^{(3)} x_{j_1}^2 + 65 m^{(4)} x_{j_1}^3 + 15 m^{(5)} x_{j_1}^4 + m^{(6)} x_{j_1}^5\big], \\[1.1mm]
        \EE[\xi_{j_1}^5 \xi_{j_2}]
        &= x_{j_1} x_{j_2} \big[m^{(2)} + 15 m^{(3)} x_{j_1} + 25 m^{(4)} x_{j_1}^2 + 10 m^{(5)} x_{j_1}^3 + m^{(6)} x_{j_1}^4\big], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}^2]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + m^{(3)} (7 x_{j_1} + x_{j_2}) + m^{(4)} (6 x_{j_1}^2 + 7 x_{j_1} x_{j_2}) \\[0.49mm]
                + m^{(5)} (x_{j_1}^3 + 6 x_{j_1}^2 x_{j_2}) + m^{(6)} x_{j_1}^3 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2} \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \big[m^{(3)} + 7 m^{(4)} x_{j_1} + 6 m^{(5)} x_{j_1}^2 + m^{(6)} x_{j_1}^3\big], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^3]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + m^{(3)} (3 x_{j_1} + 3 x_{j_2}) + m^{(4)} (x_{j_1}^2 + 9 x_{j_1} x_{j_2} + x_{j_2}^2) \\[0.49mm]
                + m^{(5)} (3 x_{j_1}^2 x_{j_2} + 3 x_{j_1} x_{j_2}^2) + m^{(6)} x_{j_1}^2 x_{j_2}^2
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^2 \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \big[m^{(3)} + m^{(4)} (3 x_{j_1} + x_{j_2}) + m^{(5)} (x_{j_1}^2 + 3 x_{j_1} x_{j_2}) + m^{(6)} x_{j_1}^2 x_{j_2}\big], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2} \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \big[m^{(4)} + 3 m^{(5)} x_{j_1} + m^{(6)} x_{j_1}^2\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3}^2]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (x_{j_1} + x_{j_2} + x_{j_3}) \\[0.49mm]
                + m^{(5)} (x_{j_1} x_{j_2} + x_{j_1} x_{j_3} + x_{j_2} x_{j_3}) + m^{(6)} x_{j_1} x_{j_2} x_{j_3}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \big[m^{(4)} + m^{(5)} (x_{j_1} + x_{j_2}) + m^{(6)} x_{j_1} x_{j_2}\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} \big[m^{(5)} + m^{(6)} x_{j_1}\big], \\[1.1mm]
        \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5} \xi_{j_6}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} x_{j_6} m^{(6)}.
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $7$:} For any distinct $j_1, j_2, j_3, j_4, j_5, j_6, j_7\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}^7]
        &= x_{j_1} \left[\hspace{-1mm}
            \begin{array}{l}
                m + 63 m^{(2)} x_{j_1} + 301 m^{(3)} x_{j_1}^2 + 350 m^{(4)} x_{j_1}^3 \\[0.49mm]
                + 140 m^{(5)} x_{j_1}^4 + 21 m^{(6)} x_{j_1}^5 + m^{(7)} x_{j_1}^6
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^6 \xi_{j_2}]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + 31 m^{(3)} x_{j_1} + 90 m^{(4)} x_{j_1}^2 \\[0.49mm]
                + 65 m^{(5)} x_{j_1}^3 + 15 m^{(6)} x_{j_1}^4 + m^{(7)} x_{j_1}^5
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^5 \xi_{j_2}^2]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + m^{(3)} (15 x_{j_1} + x_{j_2}) + m^{(4)} (25 x_{j_1}^2 + 15 x_{j_1} x_{j_2}) \\[0.49mm]
                + m^{(5)} (10 x_{j_1}^3 + 25 x_{j_1}^2 x_{j_2}) + m^{(6)} (x_{j_1}^4 + 10 x_{j_1}^3 x_{j_2}) + m^{(7)} x_{j_1}^4 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^5 \xi_{j_2} \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \big[m^{(3)} + 15 m^{(4)} x_{j_1} + 25 m^{(5)} x_{j_1}^2 + 10 m^{(6)} x_{j_1}^3 + m^{(7)} x_{j_1}^4\big], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}^3]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + m^{(3)} (7 x_{j_1} + 3 x_{j_2}) + m^{(4)} (6 x_{j_1}^2 + 21 x_{j_1} x_{j_2} + x_{j_2}^2) \\[0.49mm]
                + m^{(5)} (x_{j_1}^3 + 18 x_{j_1}^2 x_{j_2} + 7 x_{j_1} x_{j_2}^2) \\[0.49mm]
                + m^{(6)} (3 x_{j_1}^3 x_{j_2} + 6 x_{j_1}^2 x_{j_2}^2) + m^{(7)} x_{j_1}^3 x_{j_2}^2
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}^2 \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (7 x_{j_1} + x_{j_2}) + m^{(5)} (6 x_{j_1}^2 + 7 x_{j_1} x_{j_2}) \\[0.49mm]
                + m^{(6)} (x_{j_1}^3 + 6 x_{j_1}^2 x_{j_2}) + m^{(7)} x_{j_1}^3 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2} \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \big[m^{(4)} + 7 m^{(5)} x_{j_1} + 6 m^{(6)} x_{j_1}^2 + m^{(7)} x_{j_1}^3\big], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^3 \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (3 x_{j_1} + 3 x_{j_2}) + m^{(5)} (x_{j_1}^2 + 9 x_{j_1} x_{j_2} + x_{j_2}^2) \\[0.49mm]
                + m^{(6)} (3 x_{j_1}^2 x_{j_2} + 3 x_{j_1} x_{j_2}^2) + m^{(7)} x_{j_1}^2 x_{j_2}^2
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^2 \xi_{j_3}^2]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (3 x_{j_1} + x_{j_2} + x_{j_3}) \\[0.49mm]
                + m^{(5)} (x_{j_1}^2 + 3 x_{j_1} x_{j_2} + 3 x_{j_1} x_{j_3} + x_{j_2} x_{j_3}) \\[0.49mm]
                + m^{(6)} (x_{j_1}^2 x_{j_2} + x_{j_1}^2 x_{j_3} + 3 x_{j_1} x_{j_2} x_{j_3}) + m^{(7)} x_{j_1}^2 x_{j_2} x_{j_3}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^2 \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(4)} + m^{(5)} (3 x_{j_1} + x_{j_2}) \\[0.49mm]
                + m^{(6)} (x_{j_1}^2 + 3 x_{j_1} x_{j_2}) + m^{(7)} x_{j_1}^2 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} \big[m^{(5)} + 3 m^{(6)} x_{j_1} + m^{(7)} x_{j_1}^2\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3}^2 \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(4)} + m^{(5)} (x_{j_1} + x_{j_2} + x_{j_3}) \\[0.49mm]
                + m^{(6)} (x_{j_1} x_{j_2} + x_{j_1} x_{j_3} + x_{j_2} x_{j_3}) + m^{(7)} x_{j_1} x_{j_2} x_{j_3}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3} \xi_{j_4} \xi_{j_5}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} \big[m^{(5)} + m^{(6)} (x_{j_1} + x_{j_2}) + m^{(7)} x_{j_1} x_{j_2}\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5} \xi_{j_6}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} x_{j_6} \big[m^{(6)} + m^{(7)} x_{j_1}\big], \\[1.1mm]
        \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5} \xi_{j_6} \xi_{j_7}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} x_{j_6} x_{j_7} m^{(7)}.
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $8$:} For any distinct $j_1, j_2, j_3, j_4, j_5, j_6, j_7, j_8\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1}^8]
        &= x_{j_1} \left[\hspace{-1mm}
            \begin{array}{l}
                m + 127 m^{(2)} x_{j_1} + 966 m^{(3)} x_{j_2}^2 + 1701 m^{(4)} x_{j_1}^3 \\[0.49mm]
                + 1050 m^{(5)} x_{j_1}^4 + 266 m^{(6)} x_{j_1}^5 + 28 m^{(7)} x_{j_1}^6 + m^{(8)} x_{j_1}^7
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^7 \xi_{j_2}]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m + 63 m^{(3)} x_{j_1} + 301 m^{(4)} x_{j_1}^2 + 350 m^{(5)} x_{j_1}^3 \\[0.49mm]
                + 140 m^{(6)} x_{j_1}^4 + 21 m^{(7)} x_{j_1}^5 + m^{(8)} x_{j_1}^6
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^6 \xi_{j_2}^2]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + m^{(3)} (31 x_{j_1} + x_{j_2}) + m^{(4)} (90 x_{j_1}^2 + 31 x_{j_1} x_{j_2}) \\[0.49mm]
                + m^{(5)} (65 x_{j_1}^3 + 90 x_{j_1}^2 x_{j_2}) + m^{(6)} (15 x_{j_1}^4 + 65 x_{j_1}^3 x_{j_2}) \\[0.49mm]
                + m^{(7)} (x_{j_1}^5 + 15 x_{j_1}^4 x_{j_2}) + m^{(8)} x_{j_1}^5 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^6 \xi_{j_2} \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + 31 m^{(4)} x_{j_1} + 90 m^{(5)} x_{j_1}^2 \\[0.49mm]
                + 65 m^{(6)} x_{j_1}^3 + 15 m^{(7)} x_{j_1}^4 + m^{(8)} x_{j_1}^5
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^5 \xi_{j_2}^3]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + m^{(3)} (15 x_{j_1} + 3 x_{j_2}) + m^{(4)} (25 x_{j_1}^2 + 45 x_{j_1} x_{j_2} + x_{j_2}^2) \\[0.49mm]
                + m^{(5)} (10 x_{j_1}^3 + 75 x_{j_1}^2 x_{j_2} + 15 x_{j_1} x_{j_2}^2) \\[0.49mm]
                + m^{(6)} (x_{j_4}^4 + 30 x_{j_1}^3 x_{j_2} + 25 x_{j_1}^2 x_{j_2}^2) \\[0.49mm]
                + m^{(7)} (3 x_{j_1}^4 x_{j_2} + 10 x_{j_1}^3 x_{j_2}^2) + m^{(8)} x_{j_1}^4 x_{j_2}^2
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^5 \xi_{j_2}^2 \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (15 x_{j_1} + x_{j_2}) + m^{(5)} (25 x_{j_1}^2 + 15 x_{j_1} x_{j_2}) \\[0.49mm]
                + m^{(6)} (10 x_{j_1}^3 + 25 x_{j_1}^2 x_{j_2}) + m^{(7)} (x_{j_1}^4 + 10 x_{j_1}^3 x_{j_2}) + m^{(8)} x_{j_1}^4 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^5 \xi_{j_2} \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \big[m^{(4)} + 15 m^{(5)} x_{j_1} + 25 m^{(6)} x_{j_1}^2 + 10 m^{(7)} x_{j_1}^3 + m^{(8)} x_{j_1}^4\big], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}^4]
        &= x_{j_1} x_{j_2} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(2)} + m^{(3)} (7 x_{j_1} + 7 x_{j_2}) + m^{(4)} (6 x_{j_1}^2 + 49 x_{j_1} x_{j_2} + 6 x_{j_2}^2) \\[0.49mm]
                + m^{(5)} (x_{j_1}^3 + 42 x_{j_1}^2 x_{j_2} + 42 x_{j_1} x_{j_2}^2 + x_{j_2}^3) \\[0.49mm]
                + m^{(6)} (7 x_{j_1}^3 x_{j_2} + 36 x_{j_1}^2 x_{j_2}^2 + 7 x_{j_1} x_{j_2}^3) \\[0.49mm]
                + m^{(7)} (6 x_{j_1}^3 x_{j_2}^2 + 6 x_{j_1}^2 x_{j_2}^3) + m^{(8)} x_{j_1}^3 x_{j_2}^3
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}^3 \xi_{j_3}]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (7 x_{j_1} + 3 x_{j_2}) + m^{(5)} (6 x_{j_1}^2 + 21 x_{j_1} x_{j_2} + x_{j_2}^2) \\[0.49mm]
                + m^{(6)} (x_{j_1}^3 + 18 x_{j_1}^2 x_{j_2} + 7 x_{j_1} x_{j_2}^2) \\[0.49mm]
                + m^{(7)} (3 x_{j_1}^3 x_{j_2} + 6 x_{j_1}^2 x_{j_2}^2) + m^{(8)} x_{j_1}^3 x_{j_2}^2
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}^2 \xi_{j_3}^2]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (7 x_{j_1} + x_{j_2} + x_{j_3}) \\[0.49mm]
                + m^{(5)} (6 x_{j_1}^2 + 7 x_{j_1} x_{j_2} + 7 x_{j_1} x_{j_3} + x_{j_2} x_{j_3}) \\[0.49mm]
                + m^{(6)} (x_{j_1}^3 + 6 x_{j_1}^2 x_{j_2} + 6 x_{j_1}^2 x_{j_3} + 7 x_{j_1} x_{j_2} x_{j_3}) \\[0.49mm]
                + m^{(7)} (x_{j_1}^3 x_{j_2} + x_{j_1}^3 x_{j_3} + 6 x_{j_1}^2 x_{j_2} x_{j_3}) + m^{(8)} x_{j_1}^3 x_{j_2} x_{j_3}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2}^2 \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(4)} + m^{(5)} (7 x_{j_1} + x_{j_2}) + m^{(6)} (6 x_{j_1}^2 + 7 x_{j_1} x_{j_2}) \\[0.49mm]
                + m^{(7)} (x_{j_1}^3 + 6 x_{j_1}^2 x_{j_2}) + m^{(8)} x_{j_1}^3 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^4 \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} \big[m^{(5)} + 7 m^{(6)} x_{j_1} + 6 m^{(7)} x_{j_1}^2 + m^{(8)} x_{j_1}^3\big], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^3 \xi_{j_3}^2]
        &= x_{j_1} x_{j_2} x_{j_3} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(3)} + m^{(4)} (3 x_{j_1} + 3 x_{j_2} + x_{j_3}) \\[0.49mm]
                + m^{(5)} (x_{j_1}^2 + x_{j_2}^2 + 3 x_{j_1} x_{j_3} + 3 x_{j_2} x_{j_3} + 9 x_{j_1} x_{j_2}) \\[0.49mm]
                + m^{(6)} (x_{j_1}^2 x_{j_3} + x_{j_2}^2 x_{j_3} + 3 x_{j_1}^2 x_{j_2} + 3 x_{j_1} x_{j_2}^2 + 9 x_{j_1} x_{j_2} x_{j_3}) \\[0.49mm]
                + m^{(7)} (x_{j_1}^2 x_{j_2}^2 + 3 x_{j_1}^2 x_{j_2} x_{j_3} + 3 x_{j_1} x_{j_2}^2 x_{j_3}) + m^{(8)} x_{j_1}^2 x_{j_2}^2 x_{j_3}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^3 \xi_{j_3} \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(4)} + m^{(5)} (3 x_{j_1} + 3 x_{j_2}) + m^{(6)} (x_{j_1}^2 + 9 x_{j_1} x_{j_2} + x_{j_2}^2) \\[0.49mm]
                + m^{(7)} (3 x_{j_1}^2 x_{j_2} + 3 x_{j_1} x_{j_2}^2) + m^{(8)} x_{j_1}^2 x_{j_2}^2
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^2 \xi_{j_3}^2 \xi_{j_4}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(4)} + m^{(5)} (3 x_{j_1} + x_{j_2} + x_{j_3}) \\[0.49mm]
                + m^{(6)} (3 x_{j_1} x_{j_2} + 3 x_{j_1} x_{j_3} + x_{j_2} x_{j_3}) \\[0.49mm]
                + m^{(7)} (x_{j_1}^2 x_{j_2} + x_{j_1}^2 x_{j_3} + 3 x_{j_1} x_{j_2} x_{j_3}) + m^{(8)} x_{j_1}^2 x_{j_2} x_{j_3}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2}^2 \xi_{j_3} \xi_{j_4} \xi_{j_5}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(5)} + m^{(6)} (3 x_{j_1} + x_{j_2}) \\[0.49mm]
                + m^{(7)} (x_{j_1}^2 + 3 x_{j_1} x_{j_2}) + m^{(8)} x_{j_1}^2 x_{j_2}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^3 \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5} \xi_{j_6}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} x_{j_6} \big[m^{(6)} + 3 m^{(7)} x_{j_1} + m^{(8)} x_{j_1}^2\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3}^2 \xi_{j_4}^2]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(4)} + m^{(5)} (x_{j_1} + x_{j_2} + x_{j_3} + x_{j_4}) \\[0.49mm]
                + m^{(6)} (x_{j_1} x_{j_2} + x_{j_1} x_{j_3} + x_{j_1} x_{j_4} + x_{j_2} x_{j_3} + x_{j_2} x_{j_4} + x_{j_3} x_{j_4}) \\[0.49mm]
                + m^{(7)} (x_{j_1} x_{j_2} x_{j_3} + x_{j_1} x_{j_2} x_{j_4} + x_{j_1} x_{j_3} x_{j_4} + x_{j_2} x_{j_3} x_{j_4}) \\[0.49mm]
                + m^{(8)} x_{j_1} x_{j_2} x_{j_3} x_{j_4}
            \end{array}
            \hspace{-1mm}\right], \\[-3.5mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3}^2 \xi_{j_4} \xi_{j_5}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} \left[\hspace{-1mm}
            \begin{array}{l}
                m^{(5)} + m^{(6)} (x_{j_1} + x_{j_2} + x_{j_3}) \\[0.49mm]
                + m^{(7)} (x_{j_1} x_{j_2} + x_{j_1} x_{j_3} + x_{j_2} x_{j_3}) + m^{(8)} x_{j_1} x_{j_2} x_{j_3}
            \end{array}
            \hspace{-1mm}\right], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2}^2 \xi_{j_3} \xi_{j_4} \xi_{j_5} \xi_{j_6}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} x_{j_6} \big[m^{(6)} + m^{(7)} (x_{j_1} + x_{j_2}) + m^{(8)} x_{j_1} x_{j_2}\big], \\[1.1mm]
        \EE[\xi_{j_1}^2 \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5} \xi_{j_6} \xi_{j_7}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} x_{j_6} x_{j_7} \big[m^{(7)} + m^{(8)} x_{j_1}\big], \\[1.1mm]
        \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3} \xi_{j_4} \xi_{j_5} \xi_{j_6} \xi_{j_7} \xi_{j_8}]
        &= x_{j_1} x_{j_2} x_{j_3} x_{j_4} x_{j_5} x_{j_6} x_{j_7} x_{j_8} m^{(8)}.
    \end{align}

    \subsection{Computation of the central moments up to order~4}

    With the results of the previous subsection and some algebraic manipulations (or the formula in Theorem~\ref{thm:central.moments}), we can now calculate the central moments explicitly.
    We could calculate them up to order~8, but it would be very tedious.
    Instead, we write them up to order~4 for the sake of brevity.
    The simplifications we make to obtain the boxed expressions below are done with \texttt{Mathematica}.

    \iffalse
    \noindent
    \underline{Order $1$:} For any $j_1\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[\xi_{j_1} - \EE[\xi_{j_1}]] = \boxed{0}\, .
    \end{align}
    \fi

    \vspace{3mm}
    \noindent
    \underline{Order $2$:} For any distinct $j_1, j_2\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[(\xi_{j_1} - \EE[\xi_{j_1}])^2]
        &= \EE[\xi_{j_1}^2] - (\EE[\xi_{j_1}])^2
        = x_{j_1} \big[m + m^{(2)} x_{j_1}\big] - m^2 x_{j_1}^2 \notag \\[1.1mm]
        &= \boxed{m x_{j_1} (1 - x_{j_1})} \\[3.1mm]
        \EE[(\xi_{j_1} - \EE[\xi_{j_1}]) (\xi_{j_2} - \EE[\xi_{j_2}])]
        &= \EE[\xi_{j_1} \xi_{j_2}] - \EE[\xi_{j_1}] \EE[\xi_{j_2}]
        = m^{(2)} x_{j_1} x_{j_2} - m x_{j_1} m x_{j_2} \notag \\[1.1mm]
        &= \boxed{- m x_{j_1} x_{j_2}} \, .
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $3$:} For any distinct $j_1, j_2, j_3\in \{1,2,\dots,d\}$,
    \begin{align}
        \EE[(\xi_{j_1} - \EE[\xi_{j_1}])^3]
        &= \EE[\xi_{j_1}^3] - 3 \, \EE[\xi_{j_1}^2] \EE[\xi_{j_1}] + 2 \, (\EE[\xi_{j_1}])^3 \notag \\[1.1mm]
        &= x_{j_1} \big[m + 3 m^{(2)} x_{j_1} + m^{(3)} x_{j_1}^2\big] - 3 x_{j_1} \big[m + m^{(2)} x_{j_1}\big] m x_{j_1} + 2 m^3 x_{j_1}^3 \notag \\[1.1mm]
        &= \boxed{m x_{j_1} (x_{j_1} - 1) (2 x_{j_1} - 1)} \\[3.1mm]
        &\hspace{-23.6mm}\EE[(\xi_{j_1} - \EE[\xi_{j_1}])^2 (\xi_{j_2} - \EE[\xi_{j_2}])] \notag \\[1.1mm]
        &= \EE[\xi_{j_1}^2 \xi_{j_2}] - \EE[\xi_{j_1}^2] \EE[\xi_{j_2}] - 2 \, \EE[\xi_{j_1} \xi_{j_2}] \EE[\xi_{j_1}] + 2 \, (\EE[\xi_{j_1}])^2 \EE[\xi_{j_2}] \notag \\[1.1mm]
        &= x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} x_{j_1}\big] - x_{j_1} \big[m + m^{(2)} x_{j_1}\big] m x_{j_2} - 2 m^{(2)} x_{j_1} x_{j_2} m x_{j_1} + 2 m^2 x_{j_1}^2 m x_{j_2} \notag \\[1.1mm]
        &= \boxed{m x_{j_1} x_{j_2} (2 x_{j_1} - 1)} \\[3.1mm]
        &\hspace{-23.6mm}\EE[(\xi_{j_1} - \EE[\xi_{j_1}]) (\xi_{j_2} - \EE[\xi_{j_2}]) (\xi_{j_3} - \EE[\xi_{j_3}])] \notag \\[1.1mm]
        &= \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3}] - \EE[\xi_{j_1} \xi_{j_2}] \EE[\xi_{j_3}] - \EE[\xi_{j_1} \xi_{j_3}] \EE[\xi_{j_2}] - \EE[\xi_{j_2} \xi_{j_3}] \EE[\xi_{j_1}] + 2 \, \EE[\xi_{j_1}] \EE[\xi_{j_2}] \EE[\xi_{j_3}] \notag \\[1.1mm]
        &= m^{(3)} x_{j_1} x_{j_2} x_{j_3} - m^{(2)} x_{j_1} x_{j_2} m x_{j_3} - m^{(2)} x_{j_1} x_{j_3} m x_{j_2} - m^{(2)} x_{j_2} x_{j_3} m x_{j_1} + 2 m^3 x_{j_1} x_{j_2} x_{j_3} \notag \\[1.1mm]
        &= \boxed{2 m x_{j_1} x_{j_2} x_{j_3}} \, .
    \end{align}

    \vspace{2mm}
    \noindent
    \underline{Order $4$:} For any distinct $j_1, j_2, j_3, j_4\in \{1,2,\dots,d\}$,
    \begin{align}
        &\EE[(\xi_{j_1} - \EE[\xi_{j_1}])^4] \notag \\[1.1mm]
        &\qquad= \EE[\xi_{j_1}^4] - 4 \, \EE[\xi_{j_1}^3] \EE[\xi_{j_1}] + 6 \, \EE[\xi_{j_1}^2] (\EE[\xi_{j_1}])^2 - 3 \, (\EE[\xi_{j_1}])^4 \notag \\[1.1mm]
        &\qquad= x_{j_1} \big[m + 7 m^{(2)} x_{j_1} + 6 m^{(3)} x_{j_1}^2 + m^{(4)} x_{j_1}^3\big] - 4 x_{j_1} \big[m + 3 m^{(2)} x_{j_1} + m^{(3)} x_{j_1}^2\big] m x_{j_1} \notag \\
        &\qquad\quad+ 6 x_{j_1} \big[m + m^{(2)} x_{j_1}\big] (m x_{j_1})^2 - 3 m^4 x_{j_1}^4 \notag \\
        &\qquad= \boxed{3 m^2 x_{j_1}^2 (x_{j_1} - 1)^2 + m x_{j_1} (1 - x_{j_1}) (6 x_{j_1}^2 - 6 x_{j_1} + 1)} \\[2.1mm]
        %%%
        &\EE[(\xi_{j_1} - \EE[\xi_{j_1}])^3 (\xi_{j_2} - \EE[\xi_{j_2}])] \notag \\[1.1mm]
        &\qquad= \EE[\xi_{j_1}^3 \xi_{j_2}] - \EE[\xi_{j_1}^3] \EE[\xi_{j_2}] - 3 \, \EE[\xi_{j_1}^2 \xi_{j_2}] \EE[\xi_{j_1}] + 3 \, \EE[\xi_{j_1}^2] \EE[\xi_{j_1}] \EE[\xi_{j_2}] \notag \\[1.1mm]
        &\qquad\quad+ 3 \, \EE[\xi_{j_1} \xi_{j_2}] (\EE[\xi_{j_1}])^2 - 3 \, (\EE[\xi_{j_1}])^3 \EE[\xi_{j_2}] \notag \\[1.1mm]
        &\qquad= x_{j_1} x_{j_2} \big[m^{(2)} + 3 m^{(3)} x_{j_1} + m^{(4)} x_{j_1}^2\big] - x_{j_1} \big[m + 3 m^{(2)} x_{j_1} + m^{(3)} x_{j_1}^2\big] m x_{j_2} \notag \\
        &\qquad\quad- 3 x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} x_{j_1}\big] m x_{j_1} + 3 x_{j_1} \big[m + m^{(2)} x_{j_1}\big] m x_{j_1} m x_{j_2} + 3 m^{(2)} x_{j_1} x_{j_2} m^2 x_{j_1}^2 - 3 m^3 x_{j_1}^3 m x_{j_2} \notag \\
        &\qquad= \boxed{m x_{j_1} x_{j_2} (3 (m - 2) x_{j_1} (x_{j_1} - 1) - 1)} \\[2.1mm]
        %%%
        &\EE[(\xi_{j_1} - \EE[\xi_{j_1}])^2 (\xi_{j_2} - \EE[\xi_{j_2}])^2] \notag \\[1.1mm]
        &\qquad= \EE[\xi_{j_1}^2 \xi_{j_2}^2] - 2 \, \EE[\xi_{j_1}^2 \xi_{j_2}] \EE[\xi_{j_2}] - 2 \, \EE[\xi_{j_1} \xi_{j_2}^2] \EE[\xi_{j_1}] + \EE[\xi_{j_1}^2] (\EE[\xi_{j_2}])^2 + \EE[\xi_{j_2}^2] (\EE[\xi_{j_1}])^2 \notag \\[1.1mm]
        &\qquad\quad+ 4 \, \EE[\xi_{j_1} \xi_{j_2}] \EE[\xi_{j_1}] \EE[\xi_{j_2}] - 3 \, (\EE[\xi_{j_1}])^2 (\EE[\xi_{j_2}])^2 \notag \\[1.1mm]
        &\qquad= x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} (x_{j_1} + x_{j_2}) + m^{(4)} x_{j_1} x_{j_2}\big] - 2 x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} x_{j_1}\big] m x_{j_2} \notag \\
        &\qquad\quad- 2 x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} x_{j_2}\big] m x_{j_1} + x_{j_1} \big[m + m^{(2)} x_{j_1}\big] m^2 x_{j_2}^2 + x_{j_2} \big[m + m^{(2)} x_{j_2}\big] m^2 x_{j_1}^2 \notag \\
        &\qquad\quad+ 4 m^{(2)} x_{j_1} x_{j_2} m x_{j_1} m x_{j_2} - 3 \, m^2 x_{j_1}^2 m^2 x_{j_2}^2 \notag \\
        &\qquad= \boxed{m (m - 2) x_{j_1} x_{j_2} (3 x_{j_1} x_{j_2} - (x_{j_1} + x_{j_2}) + 1) + m x_{j_1} x_{j_2}} \\[2.1mm]
        %%%
        &\EE[(\xi_{j_1} - \EE[\xi_{j_1}])^2 (\xi_{j_2} - \EE[\xi_{j_2}]) (\xi_{j_3} - \EE[\xi_{j_3}])] \notag \\[1.1mm]
        &\qquad= \EE[\xi_{j_1}^2 \xi_{j_2} \xi_{j_3}] - \EE[\xi_{j_1}^2 \xi_{j_2}] \EE[\xi_{j_3}] - \EE[\xi_{j_1}^2 \xi_{j_3}] \EE[\xi_{j_2}] - 2 \, \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3}] \EE[\xi_{j_1}] + \EE[\xi_{j_1}^2] \EE[\xi_{j_2}] \EE[\xi_{j_3}] \notag \\[1.1mm]
        &\qquad\quad+ 2 \, \EE[\xi_{j_1} \xi_{j_2}] \EE[\xi_{j_1}] \EE[\xi_{j_3}] + 2 \, \EE[\xi_{j_1} \xi_{j_3}] \EE[\xi_{j_1}] \EE[\xi_{j_2}] + \EE[\xi_{j_2} \xi_{j_3}] (\EE[\xi_{j_1}])^2 - 3 \, (\EE[\xi_{j_1}])^2 \EE[\xi_{j_2}] \EE[\xi_{j_3}] \notag \\[1.1mm]
        &\qquad= x_{j_1} x_{j_2} x_{j_3} \big[m^{(3)} + m^{(4)} x_{j_1}\big] - x_{j_1} x_{j_2} \big[m^{(2)} + m^{(3)} x_{j_1}] m x_{j_3} - x_{j_1} x_{j_3} \big[m^{(2)} + m^{(3)} x_{j_1}] m x_{j_2} \notag \\
        &\qquad\quad- 2 m^{(3)} x_{j_1} x_{j_2} x_{j_3} m x_{j_1} + x_{j_1} \big[m + m^{(2)} x_{j_1}\big] m x_{j_2} m x_{j_3} + 2 m^{(2)} x_{j_1} x_{j_2} m x_{j_1} m x_{j_3}  \notag \\
        &\qquad\quad+ 2 m^{(2)} x_{j_1} x_{j_3} m x_{j_1} m x_{j_2} + m^{(2)} x_{j_2} x_{j_3} m^2 x_{j_1}^2 - 3 m^2 x_{j_1}^2 m x_{j_2} m x_{j_3} \notag \\
        &\qquad= \boxed{m (m - 2) x_{j_1} x_{j_2} x_{j_3} (3 x_{j_1} - 1)} \\[2.1mm]
        %%%
        &\EE[(\xi_{j_1} - \EE[\xi_{j_1}]) (\xi_{j_2} - \EE[\xi_{j_2}]) (\xi_{j_3} - \EE[\xi_{j_3}]) (\xi_{j_4} - \EE[\xi_{j_4}])] \notag \\[1.1mm]
        &\qquad= \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3} \xi_{j_4}] - \EE[\xi_{j_1} \xi_{j_2} \xi_{j_3}] \EE[\xi_{j_4}] - \EE[\xi_{j_1} \xi_{j_2} \xi_{j_4}] \EE[\xi_{j_3}] - \EE[\xi_{j_1} \xi_{j_3} \xi_{j_4}] \EE[\xi_{j_2}] - \EE[\xi_{j_2} \xi_{j_3} \xi_{j_4}] \EE[\xi_{j_1}] \notag \\[1.1mm]
        &\qquad\quad+ \EE[\xi_{j_1} \xi_{j_2}] \EE[\xi_{j_3}] \EE[\xi_{j_4}] + \EE[\xi_{j_1} \xi_{j_3}] \EE[\xi_{j_2}] \EE[\xi_{j_4}] + \EE[\xi_{j_1} \xi_{j_4}] \EE[\xi_{j_2}] \EE[\xi_{j_3}] \notag \\[1.1mm]
        &\qquad\quad+ \EE[\xi_{j_2} \xi_{j_3}] \EE[\xi_{j_1}] \EE[\xi_{j_4}] + \EE[\xi_{j_2} \xi_{j_4}] \EE[\xi_{j_1}] \EE[\xi_{j_3}] + \EE[\xi_{j_3} \xi_{j_4}] \EE[\xi_{j_1}] \EE[\xi_{j_2}] - 3\, \EE[\xi_{j_1}] \EE[\xi_{j_2}] \EE[\xi_{j_3}] \EE[\xi_{j_4}] \notag \\[1.1mm]
        &\qquad= m^{(4)} x_{j_1} x_{j_2} x_{j_3} x_{j_4} - m^{(3)} x_{j_1} x_{j_2} x_{j_3} m x_{j_4} - m^{(3)} x_{j_1} x_{j_2} x_{j_4} m x_{j_3} - m^{(3)} x_{j_1} x_{j_3} x_{j_4} m x_{j_2} - m^{(3)} x_{j_2} x_{j_3} x_{j_4} m x_{j_1} \notag \\
        &\qquad\quad+ m^{(2)} x_{j_1} x_{j_2} m x_{j_3} m x_{j_4} + m^{(2)} x_{j_1} x_{j_3} m x_{j_2} m x_{j_4} + m^{(2)} x_{j_1} x_{j_4} m x_{j_2} m x_{j_3} \notag \\
        &\qquad\quad+ m^{(2)} x_{j_2} x_{j_3} m x_{j_1} m x_{j_4} + m^{(2)} x_{j_2} x_{j_4} m x_{j_1} m x_{j_3} + m^{(2)} x_{j_3} x_{j_4} m x_{j_1} m x_{j_2} - 3\, m^4 x_{j_1} x_{j_2} x_{j_3} x_{j_4} \notag \\
        &\qquad= \boxed{3 m (m - 2) x_{j_1} x_{j_2} x_{j_3} x_{j_4}}\, .
    \end{align}

\section{Conclusion}

    In this short paper, we found general formulas for the central and non-central moments of the multinomial distribution as well as explicit formulas for all the non-central moments up to order~8 and all the central moments up to order~4.
    Our work expands on the results in \cite{Newcomer2008phd}, where the central moments were calculated up to order~4.
    It also complements the general formula for the (joint) factorial moments from \cite{MR143299} and the explicit formulas for some of the lower-order (mixed) cumulants that were presented in \cite{MR33996}.

%
% ----------  B I B L I O G R A P H I E  ----------
%

\nocite{MR3825458}
\nocite{Ouimet2019phd}

\bibliographystyle{authordate1}
\bibliography{Ouimet_2020_moments_multinomial_arxiv_bib}

\end{document}